The de Broglie wavelength of a 0.0600 gram bullet traveling at the speed of 655.5 m/s is ________ m.

is the answer:

a. 1.685×10−35

b. 1.685×10−33

c. 2.606×10−35

d. 1.685×10−32

e. 1.685×10−34
Answer #1

By using formula,

λ = h/p

    where λ is wavelength,

               h is Planck constant,

    λ = h/mv

      = 6.626*10-34 / (6*10-5) *(655.5)

      = = 1.685*10-32 m